# QWENDY: Gene Regulatory Network Inference by Quadruple Covariance Matrices

### v3

@@ 5 Discussion
Although QWENDY can fully determine the GRN, the diagonal elements that represent the regulation of one gene to itself (autoregulation) can be mixed with natural mRNA synthesis and degradation [46]. Therefore, it is not recommended to infer the autoregulation with QWENDY or its variants.
@@ Acknowledgments
Appendix A Enhancing QWENDY with deep learning
A.1 Methods
@@
Inspired by TRENDY, we propose a similar solution to enhance QWENDY. See Algorithms 2,3 for details. Since $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ are only used to distinguish between $B$ and $-B$, it is not necessary to train another model for them.
@@
    Repeat generating random GRN $A_{\mathrm{true}}$ and corresponding gene expression data at time 00, $t$, $2t$, $3t$ from a gene expression simulator
@@
We name Algorithms 2,3 with TE structure as TEQWENDY, and Algorithms 2,3 with LE structure as LEQWENDY. See Appendix for details of TE and LE structures. TEQWENDY has 4.7 million trainable parameters. LEQWENDY has 4.6 million trainable parameters, with 300 million non-trainable (frozen) parameters.
@@
where $X_{i}(t)$ is the level of gene $i$ at time $t$, $W_{j}(t)$ is a standard Brownian motion, and $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$ There are $10^{5}$ training samples, each with the expression levels of 10 genes for 100 cells, measured at four time points: 0.0, 0.1, 0.2, 0.3.
@@
A.2 Performance
We also test the performance of TEQWENDY and LEQWENDY on the same data sets. See Table S1 for the scores. Compared with QWENDY, TEQWENDY performs better on synthetic data, but worse on experimental data. The overall performance of TEQWENDY is slightly better than QWENDY. Therefore, the improvement of TEQWENDY might be due to the overfitting on synthetic data, which does not work on experimental data. The performance of LEQWENDY is worse than QWENDY on both synthetic and experimental data, meaning that LLMs trained with natural language inputs might not directly help with the numerical task in GRN inference, even after task-specific fine-tuning.
@@ A.3 Technical details
    Linear embedding layer with dimension $1$ to $d=64$ ($4$ groups of $n\times n\times d$)
@@
    Linear embedding layer with dimension $d$ to $d$ ($4$ groups of $n\times n\times d$)
@@
    $7$ layers of transformer encoder with $4$ heads, dimension $d$, feedforward dimension $4d$, dropout rate $0.1$ ($4$ groups of $(n^{2})\times d$)
@@
    Linear embedding layer with dimension $d$ to $d$ ($4$ groups of $(n^{2})\times d$)
@@
    LeakyReLU activation function with $\alpha=0.1$ ($4$ groups of $n\times n\times d$)
@@
    Linear embedding layer with dimension $d$ to $1$ ($4$ groups of $(n^{2})\times 1$)
@@
The 2-D positional encoding layer incorporates spatial information of matrix to the input. It generates an $n\times n\times d$ array PE and adds it to the embedded input: For $x$ and $y$ in $1,2,\ldots,n$ and $j$ in $41,\ldots,d/$ 4,
@@
    Linear embedding layer with dimension $1$ to $d=64$ ($5$ groups of $n\times n\times d$)
@@
    $3$ layers of transformer encoder with $4$ heads, dimension $5d$, feedforward dimension $20d$, dropout rate $0.1$ ($(n^{2})\times(5d)$)
@@
    Linear embedding layer with dimension $1$ to $d=256$ ($4$ groups of $n\times n\times d$)
@@
    Linear embedding layer with dimension $d$ to $d$ ($4$ groups of $n\times n\times d$)
@@
[…]
@@
    Four different linear embedding layers with dimension $4d$ to $2d$ ($4$ groups of $(n^{2})\times(2d)$)
@@
    LeakyReLU activation function with $\alpha=0.1$ ($4$ groups of $(n^{2})\times(2d)$)
@@
    Four different linear embedding layers with dimension $2d$ to $1$ ($4$ groups of $(n^{2})\times 1$)
@@
[…]
@@
    Linear embedding layer with dimension $1$ to $d_{1}=192$ or $d_{2}=256$ ($4$ groups of $n\times n\times d_{1}$ and $1$ group of $n\times n\times d_{2}$)
@@
[…]

### v4

@@ 5 Conclusion
QWENDY performs much worse on synthetic data sets. One possibility is that gene regulation can have different dynamics, and QWENDY only works on some of them, which happen to match the experimental data sets. Another possibility is that the dynamics of generating synthetic data might not match reality.
When the gene expression is at steady state, covariance matrix is time-invariant, $K_{0}=K_{1}$. This means $L_{0}^{-1}K_{1}L_{0}^{-\mathrm{T}}=I$ with repeated eigenvalue $1$. Then condition (2) and thus Theorem 1 would fail. Therefore, QWENDY does not work at steady state. GRN inference methods like Dictys [55] should be applied instead.
@@
QWENDY directly solves the GRN, not to search for the best match in a space of possible GRNs. In comparison, WENDY is an optimization algorithm, and it can incorporate prior information about the GRN by restricting the searching space [14]. Therefore, data besides mRNA count, such as motif analysis or genetic perturbation data, cannot be used inside QWENDY. One choice is to use such data to modify the results of QWENDY. Another choice is to stop QWENDY when it limits the GRN in a finite set, and directly search for the best match for extra information.
@@
Since QWENDY is based on a linear approximation, each diagonal element of the inferred GRN that represents the effect of one gene to itself is an indistinguishable mixture of autoregulation and natural mRNA degradation [56]. Therefore, it is not recommended to infer the existence of autoregulation with QWENDY or its variants.
6 Discussion: the present and the future of GRN inference
6.1 Current status: experiments
The input of GRN inference methods is determined by biological measurement techniques. As in 2025, it is common to measure the mRNA counts of different genes in single cells. However, due to the cost of time and money, each time point generally has only hundreds of cells, much smaller than the number of genes measured. Besides, the measurement is not very accurate, that there are many zero reads, meaning the loss of corresponding mRNAs during measurement.
The major problem of the popular scRNA-seq measurement is that cells are killed, and we cannot measure one cell more than once, which makes GRN inference mathematically difficult. There are some new techniques to read mRNA count or protein count for the same cell at multiple time points and provide more information. However, they have low accuracy and only work for a small number of genes, meaning that not all genes in the GRN are guaranteed to be included.
Some measurements such as motif analysis can provide direct information about the GRN, but the information is limited. The most reliable method to obtain the GRN is gene knockout or knockdown. However, it is costly to obtain a gold standard GRN through this method. Thus there are only a few fully reliable GRNs.
On the other side, since gene expression is confined in living cells, we do not fully understand the dynamics of gene expression and regulation. Thus given a GRN, it is difficult to simulate gene expression data.
6.2 Current status: inference methods
Due to the limitation of experimental measurements, different GRNs have a chance to produce the same scRNA-seq data. Thus it is theoretically difficult to infer the GRN with high accuracy, and biologists should not have too much faith in every inferred regulation relation. Still, many methods, especially QWENDY, are significantly better than random guess and can provide more true information than false information.
The central problem for the field of GRN inference is the lack of gold standard. Only a few experimental data sets can be used for evaluation, meaning that one method with good performance might just have good luck or wisely-selected hyper-parameters. For synthetic data, although the data quantity issue is solved, we do not know whether such data reflect reality, and inference methods might overfit to the simulator.
For information-based inference methods, the advantage is that information or correlation or predictability between genes represents reliable relationship. The disadvantage is the interpretation of this relationship: direct regulation, indirect regulation, confounder, collider can all lead to information, but GRN only contains direct regulations. Besides, since one cell cannot be measured twice, it is difficult to determine which is the cause and which is the result. It is also difficult to determine whether the regulation is positive or negative.
For model-based methods, to make the problem solvable, assumptions and simplifications are necessary. For example, many model-based GRN inference methods rely on linearization and do not consider the on-off switch of genes or the count of proteins. Therefore, such models cannot fully match all experimental phenomena. This means that corresponding methods cannot reach a high accuracy.
For popular deep learning methods, they have proven their capability in various fields and could perform well in GRN inference, if there are enough data. The problem is that the amount of data required is too large, but we cannot conduct so many experiments. Besides, as shown in Appendix C, not all popular deep learning structures are suitable for GRN inference.
In sum, due to the data quality and quantity issue, there is an unsatisfactory upper bound for inference accuracy. Some well-studied data types might not have the space for a significantly better new method. Some other data types, such as the one studied in this paper, has only a few corresponding methods and deserves more research.
6.3 Future of experiments and inference methods
The future of GRN inference methods heavily depends on the type and the amount of data available, determined by the development of biotechnology.
(1) If some future research makes GRNs easy and cheap to determine through direct experiments, then the whole field of GRN inference will disappear.
(2) If it becomes easy and accurate to measure the same cell multiple times, information-based methods might have the best performance. The major obstacle of information-based methods is the difficulty to determine the direction of regulation. With real time series data, we can easily determine that the one happens earlier regulates the one happens later.
(3) If real time series data are still not applicable or reliable, but there are many scRNA-seq data sets with multiple time points, sufficiently many cells, and high accuracy, then inference methods based on models for covariance matrices (such as QWENDY) can be promising. If we can know better of gene expression and regulation, but not enough for reliable gene expression data generators, then model-based methods should be further developed.
(4) If the dynamics of gene expression and regulation can be fully understood, we can obtain very reliable gene expression data generators, and deep learning methods can flourish. However, if GRN inference is needed, then it means that the gene expression dynamics for this situation is unknown, and we cannot guarantee that the dynamics matches the data generator.
From situation (1) to situation (4), the reliability of the obtained GRN should decrease.
@@ Acknowledgments
Appendix A Performance of QWENDY on full hESC data set
The hESC data set we use was processed by Matsumoto et al. [57]. The original version has 100 genes. Since the original data have many zero reads, WENDY and GENIE3 methods would fail. Therefore, in the WENDY paper [14], we only kept genes that have nonzero reads in at least 95% cells at each time point. Then only 18 genes remain. This 18-gene hESC data set was then used in the TRENDY paper [15] and the main text of this paper.
One problem is that the 18-gene hESC data set misses some essential genes in embryonic development. Therefore, we also use the 100-gene hESC data set. Since half of 16 methods (all WENDY-based and GENIE3-based) tested in the TRENDY paper [15] fail on this data set, we only apply the QWENDY method. Besides, the ground truth GRN of this data set is based on motif analysis of open chromatin, not genetic perturbation data, making it less reliable. Therefore, we focus on essential genes in embryonic development: GATA6, NANOG, EOMES, SOX2, SOX17, SMAD2, FOXH1, GATA4, POU5F1 [58]. In the inferred GRN from the 100-gene hESC data set, we calculate the overall regulation power of each gene:
$R_{i}=\sum_{j\neq i}|B_{ij}|.$
Then we rank the regulation power of each gene from high to low. Among all 100 genes, SOX17, GATA6, and GATA4 are among the top 10. The average rank of these essential genes is 30.1. Thus QWENDY correctly infers that such essential genes have strong regulation power.
Appendix B Details of mESC data set
The mESC data set measures the gene expression at five time points: 0h, 12h, 24h, 48h, 72h. Each time point has around 100 cells. We use the mESC data set with 100 genes and corresponding ground truth GRN processed by Matsumoto et al. [57]. Since the original data have many zero reads, half of 16 methods (all WENDY-based and GENIE3-based) tested in the TRENDY paper [15] fail. Therefore, we only keep genes that have nonzero reads in at least 95% cells at each time point. Now only 34 genes remain, and we can test all methods. See Table S1 for performance of 16 methods tested in the TRENDY paper.
Table S1: AUROC and AUPRC of different methods on the mESC data set
Method AUROC AUPRC
WENDY 0.4857 0.0411
TRENDY 0.4655 0.0489
nWENDY 0.4273 0.0370
bWENDY 0.4296 0.0370
GENIE3 0.5024 0.0452
tGENIE3 0.5401 0.0556
nGENIE3 0.4779 0.0424
bGENIE3 0.4730 0.0422
SINCERITIES 0.5744 0.0630
tSINCERITIES 0.4930 0.0411
nSINCERITIES 0.5896 0.0542
bSINCERITIES 0.5755 0.0517
NonlinearODEs 0.4940 0.0517
tNonlinearODEs 0.3737 0.0323
nNonlinearODEs 0.4957 0.0518
bNonlinearODEs 0.4955 0.0514
Appendix C Enhancing QWENDY with deep learning
C.1 Methods
@@
Inspired by TRENDY, we propose a similar solution to enhance QWENDY. See Algorithms 2,3 for details. Since $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ are only used to distinguish between $B$ and B-B, it is not necessary to train another model for them.
@@
    Repeat generating random GRN $A_{\mathrm{true}}$ and corresponding gene expression data at time 0, $t$, $2t$, $3t$ from a gene expression simulator
@@
We name Algorithms 2,3 with TE structure as TEQWENDY, and Algorithms 2,3 with LE structure as LEQWENDY. See the technical details section for details of TE and LE structures. TEQWENDY has 4.7 million trainable parameters. LEQWENDY has 4.6 million trainable parameters, with 300 million non-trainable (frozen) parameters.
@@
where $X_{i}(t)$ is the level of gene $i$ at time $t$, $W_{j}(t)$ is a standard Brownian motion, and $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$ . There are $10^{5}$ training samples, each with the expression levels of 10 genes for 100 cells, measured at four time points: 0.0, 0.1, 0.2, 0.3.
@@
C.2 Performance
We also test the performance of TEQWENDY and LEQWENDY on the same data sets. See Table S2 for the scores. Compared with QWENDY, TEQWENDY performs better on synthetic data, but worse on experimental data. The overall performance of TEQWENDY is worse than QWENDY. Therefore, TEQWENDY might overfit on synthetic data, which does not work on experimental data. The performance of LEQWENDY is worse than QWENDY on both synthetic and experimental data, meaning that LLMs trained with natural language inputs might not directly help with the numerical task in GRN inference, even after task-specific fine-tuning.
@@ C.3 Technical details
    Linear embedding layer with dimension $1$ to $d=64$ (44 groups of $n\times n\times d$)
@@
    Linear embedding layer with dimension $d$ to $d$ (44 groups of $n\times n\times d$)
@@
    77 layers of transformer encoder with $4$ heads, dimension $d$, feedforward dimension $4d$, dropout rate $0.1$ (44 groups of $(n^{2})\times d$)
@@
    Linear embedding layer with dimension $d$ to $d$ (44 groups of $(n^{2})\times d$)
@@
    LeakyReLU activation function with $\alpha=0.1$ (44 groups of $n\times n\times d$)
@@
    Linear embedding layer with dimension $d$ to 11 (44 groups of $(n^{2})\times 1$)
@@
The 2-D positional encoding layer incorporates spatial information of matrix to the input. It generates an $n\times n\times d$ array PE and adds it to the embedded input: For $x$ and $y$ in $1,2,\ldots,n$ and $j$ in $1,\ldots,d/4$,
@@
    Linear embedding layer with dimension $1$ to $d=64$ (55 groups of $n\times n\times d$)
@@
    33 layers of transformer encoder with $4$ heads, dimension $5d$, feedforward dimension $20d$, dropout rate $0.1$ ($(n^{2})\times(5d)$)
@@
    Linear embedding layer with dimension $1$ to $d=256$ (44 groups of $n\times n\times d$)
@@
    Linear embedding layer with dimension $d$ to $d$ (44 groups of $n\times n\times d$)
@@
    Trainable LoRA layers with rank $r=8$ and $\text{LoRA-}\alpha=16$, added to each transformer encoder layer ($(n^{2})\times(4d)$)
@@
    Four different linear embedding layers with dimension $4d$ to $2d$ (44 groups of $(n^{2})\times(2d)$)
@@
    LeakyReLU activation function with $\alpha=0.1$ (44 groups of $(n^{2})\times(2d)$)
@@
    Four different linear embedding layers with dimension $2d$ to 11 (44 groups of $(n^{2})\times 1$)
@@
For each large pre-trained weight matrix $W$ with size $p\times q$, LoRA freezes $W$ and replace it by $W+\Delta W$. Here $\Delta W=(\text{LoRA-}\alpha/r)AB$, where the trainable $A$ has size $p\times r$, and the trainable $B$ has size $r\times q$. The total number of trainable parameters decreases from $pq$ to $(p+q)r$, since $r\ll p,q$. The scaling factor $\text{LoRA-}\alpha$ controls the update rate.
@@
    Linear embedding layer with dimension $1$ to $d_{1}=192$ or $d_{2}=256$ (44 groups of $n\times n\times d_{1}$ and $1$ group of $n\times n\times d_{2}$)
@@
    Trainable LoRA layers with rank $r=16$ and $\text{LoRA-}\alpha=32$, added to each transformer encoder layer ($(n^{2})\times d$)